Formula Sheet (Phvsire Which one of the following quantities can be determined from the slope of a velocity versus time graph for an object in motion? 20. A) distance traveled B) position C) velocity D) displacement E) acceleration A projectile is fired at an angle of 35.0° above the horizontal with an initial speed of 55.0 m/s. How long does it take the projectile to reach the highest point in its tra 21. jectory? A) 2.4s B) 9.8 s C) 3.2 s D) 1.5 s E) 4.0s The work done by a force that varies in magnitude can be obtained from which of the following choices? 22. A) B) C) D) E) by finding the area under a force versus velocity graph by calculating the average force and multiplying by the speed by finding the area under a force versus displacement graph by calculating the average force and multiplying by the displacement by either choice (c) or choice (d) A particle is in uniform circular motion about the origin of an xy coordinate system, moving clockwise with a period of 7.00 s. At one instant, its position vector (measured from the origin) is: 2-2j.Calculate the speed of the particle. 23. A) 7.00 N.s B) 2.54 m/s C) 2.10 m/s D) 1.27 m/s E) 2? m/s
Answer #1

Q20.

slope of a velocity vs time graph=change in velocity/change in time

=rate of change of velocity

=acceleration

hence option E is correct.

Q21.

initial speed=55 m/s

angle=35 degrees

time taken to reach highest point=initial speed*sin(angle)/g

=55*sin(35)/9.8

=3.22 seconds

hence option C is correct.

Q22.

work done=integration of force and displacement

hence work done is area under force vs displacement graph

hence option C is correct.

Q23.

radius=sqrt(2^2+2^2)=2.8284 m

time period=perimeter/speed

==>7=2*pi*2.8284/speed

==>speed=2.5388 m/s

hence option B is correct.